# Matryoshka: Non-Exclusive Memory Tiering via Transactional Page Migration

### v1

@@ 4.2 Real-world Applications
We continued the evaluation of Matryoshka using three representative real-world applications with unique memory access patterns: Redis [10], PageRank [9], and Liblinear [5]. We ran these three applications on top of three different platforms (Table 1) to reveal the benefits and potential limitations of current page management systems for emerging tiered memory.
@@
Key-value store: We first conducted experiments on a latency-sensitive key-value database, Redis [10]. The workload was generated from YCSB [11], using its update heavy workload A, with a 50/50 distribution of read and write operations. We carefully crafted three cases with different RSS and total operations. Note that the parameters of YCSB were set as default if they were not specified as follows.
Case 1: We set recordcount to 6 million and operationcount to 8 million. After pre-loading the dataset into the database, we used a customized tool to demote all Redis’ memory pages to the slow memory tier. Then, we started to run the experiment. The RSS of this case was 13GB.
Case 2: We increased the RSS by setting recordcount to 10 million and operationcount to 12 million. We used the same way to demote all the memory pages to the slow tier. The RSS of this case was 24GB.
Case 3: We increased the total operations while keeping the same RSS by setting recordcount to 10 million and operationcount to 12 million. After pre-loading the dataset, we did not demote any memory pages. The RSS was 24GB.
Figure 10 shows that, compared to TPP, Matryoshka consistently delivered superior performance for the latency-sensitive key-value store across all tested platforms in the three cases. The enhanced performance attributable to Matryoshka stems from its asynchronous, transactional page migration and page shadowing approaches, which merely block the execution of applications. Notably, Matryoshka demonstrated the highest performance in case 3. This is due to that a smaller amount of data (8 GB) was initially allocated on the slower tier in case 3, compared to the larger initial allocations (13 GB and 16 GB) in the other two cases.
@@
Graph-based computation: We further used PageRank [9], an application used to rank web pages. It involves iterative computations to determine the rank of each page, based on the link structures of the entire web. As the size of the dataset increases, the computational complexity also increases, making it both memory-intensive and compute-intensive. We used a benchmark suite [8] to generate a synthetic uniform-random graph comprising $2^{26}$ vertices, each with an average of 20 edges. The RSS in this experiment was 22 GB, indicating that the memory pages were distributed at both the local DRAM and remote CXL memory or PM.
@@
Figure 11 illustrates that there is negligible variance in performance between scenarios with page migrations (using Matryoshka and TPP) and without page migrations (no migration). The results suggest that: 1) For non-latency-sensitive applications, such as PageRank, using CXL memory (as in Platform A) can significantly expand the local DRAM capacity without adversely impacting application-level performance. 2) In such scenarios, page migration appears to be unnecessary. In addition, the application-level performance diminishes with PM (as observed in Platform C), partly due to PM’s higher latency compared to the CXL memory. These findings also reveal that the overhead associated with Matryoshka’s page migration minimally influences PageRank’s performance. Additionally, it is observed that among all evaluated scenarios, Memtis exhibits the least efficient performance.
@@
Machine learning: Our final evaluation of Matryoshka involved using the machine learning library Liblinear [5], known for its large-scale linear classification capabilities. We executed Liblinear with an L1 regularized logistic regression workload, using 16 threads with an RSS of 10 GB. Prior to each execution, we used our tool to demote all memory pages associated with the Liblinear workload to the slower memory tier.
Figure 12 illustrates that both Matryoshka and TPP significantly outperformed the “no migration” scenario on Platform B, which utilizes CXL-based memory tiering. However, this performance advantage was not evident on Platform C, featuring PM-based memory tiering, where all approaches exhibited comparable performance. Further analysis revealed a sharp difference in page promotions between the platforms: 12 million on Platform B compared to just 1 million on Platform C. This discrepancy is attributable to the differing speeds of CXL memory and PM. This observation indicates that the efficacy of page migration is closely linked to the performance characteristics of the underlying memory media.
@@
Our observations from Figure 11 and Figure 12 also reveal that the application-level performance was related to the distinct characteristics of platforms. For instance, Platform A represents a standard off-the-shelf machine, whereas Platform B is an engineering sample with undisclosed features, which may affect various applications differently. This implies that the presence of other prominent performance-related factors could potentially overshadow the effects of page migrations.
@@ 5 Discussions and Future Work
The key insight from Matryoshka’s evaluation is that page migration, especially under memory pressure, has a detrimental impact on overall application performance. While Matryoshka achieved graceful performance degradation and much higher performance than TPP, an approach based on synchronous page migration, its performance is sub-optimal compared to that without page migration. When the program’s working set exceeds the capacity of the fast tier, the most effective strategy is to access pages directly from their initial placement, completely disabling page migration. It is straightforward to detect memory thrashing, e.g., frequent and equal number of page demotions and promotions, and disable page migrations. However, estimating the working set size to resume page migration becomes challenging, as the working set now spans multiple tiers. It requires global memory tracking, which could be prohibitively expensive, to identify the hot data set that can potentially be migrated to the fast tier. We plan to extend Matryoshka to unilaterally throttle page promotions and monitor page demotions to effectively manage memory pressure on the fast tier. Note that this would require the development of a new page migration policy, which is orthogonal to the Matryoshka page migration mechanisms proposed in this work.
There exist difficult tradeoffs between page fault-based access tracking, such as TPP and Matryoshka, and hardware performance counter-based memory access sampling like Memtis. While page fault-based tracking effectively captures access recency, it can be potentially expensive and on the critical path of program execution. In comparison, hardware-based access sampling is off the critical path and captures access frequency. However, it is not responsive to workload changes and its accuracy relies critically on the sampling rate. One advantage of Matryoshka is that it is a page fault-based migration approach that is asynchronous and completely off the critical path. A potential future work is to integrate Matryoshka with hardware-based, access frequency tracking, such as Memtis, to enhance the current migration policy.
Lastly, we have observed application and system crashes on platform-C which has almost an order of magnitude more (persistent) memory on the capacity tier than the performance tier. The crashes occurred to all tested approaches including the vanilla Linux kernel, and to applications whose working set size is significantly larger than the capacity of fast memory. Intuitively, the Linux kernel should utilize tiered memory equally when under memory pressure. However, after reviewing system logs and Linux kernel source code, we found that Linux, and provably many other commercial OSes, follows a hierarchical approach in managing the performance and memory allocation of tiered memory – attempting to cache or allocate data from fast memory if at all possible and resorting to slow memory if failed. Although there are guardrails to spill data into slow memory, such as kswapd, these mechanisms are often not responsive enough to prevent an OOM error. It is worth noting that Matryoshka has robust shadow page reclamation and does not cause OOM errors due to page shadowing. This finding motivates further research on tiered memory allocation that takes into account both non-hierarchical strategies for reliability and hierarchical strategies for performance.
@@ 6 Conclusion
This paper introduces non-exclusive memory tiering as an alternative to the common exclusive memory tiering strategy, where each page is confined to either fast or slow memory. The proposed approach, implemented in Matryoshka, leverages transactional page migration and page shadowing to enhance page management in Linux. Unlike traditional page migration, Matryoshka ensures asynchronous migration and retains shadow copies of recently promoted pages. Through comprehensive evaluations, Matryoshka demonstrates up to 6x performance improvement over existing methods, addressing critical performance degradation issues in exclusive memory tiering, especially under memory pressure. The paper calls for further research in tiered memory-aware memory allocation.

### v2

@@ 4.2 Real-world Applications
We continued the evaluation of Nomad using three representative real-world applications with unique memory access patterns: Redis [redis], PageRank [pagerankwiki], and Liblinear [liblinear]. We ran these three applications on four different platforms (platforms (as shown in Table 1) with two casesconfigurations: 1) a small RSS (under 32 GB) working with all platforms and 2) a large RSS (over 32 GB) that only Platform only on platform C and D can run due to larger with large PM or CXL memory. In addition, we include results from a “no migration” baseline which disables page migrations to show whether tiered memory management is necessary.
Figure 11: Performance comparisons using Redis and YCSB between TPP, Memtis-Default, Memtis-QuickCool, Nomad, and “No no migration”.
@@
Key-value store. We first conducted experiments on a latency-sensitive key-value database, Redis [redis]. The workload was generated from YCSB [ycsb], using its update-heavy workload A, with a 50/50 distribution of read and write operations. We crafted three cases with different RSS and total operations. Note that the parameters of YCSB were set as default unless otherwise specified. Case 1: We set recordcount to 6 million and operationcount to 8 million. After pre-loading the datasetinto the database, we used a customized tool to demote all memory pages to the slow tier before starting the experiment. The RSS of this case was 13GB. Case 2: We increased the RSS by setting recordcount to 10 million and operationcount to 12 million. We demoted all the memory pages to the slow tier in the same way. The RSS of this case was 24GB. Case 3: We kept the same total operations and RSS as Case 2. However, after pre-loading the dataset, we did not demote any memory pages.
Consistent with the micro-benchmarking results, Figure 11 shows that Nomad delivered superior performance (in terms of operations per second) compared to TPP across all platforms in all cases. In addition, Nomad outperformed Memtis when the WSS was small (i.e., for Case in case 1), but suffered more performance degradation as the WSS increased (i.e., for Case in case 2 and 3) due to an increased number of page migrations and additional overhead. Finally, all the page migration approaches underperformed compared to the “no migration” scenariobaseline. It is because the memory accesses generated by the YCSB workload were mostly “random,””, rendering migrating pages to the fast tier less effective, as those pages were unlikely to be accessed againduring the test. It indicates once again that page migration could incur nontrivial overhead, and a strategy to smartly dynamically switch it on/off is needed.
We further increased the RSS of the database and operations of YCSB by setting the recordcount to 20 million and operationcount to 30 million. The RSS for this case was 36.5GB, exceeding the total size of the tiered memory on Platforms platforms A and B, thus only being able to run on Platforms . Thus, the large RSS test was only performed on platforms C and D. We tested two initial memory placement strategies for the database – 1) Thrashedthrashing (demoting all memory pages that allocated all pages first to the slow tier ) and and immediately invoked intensive page migrations, and 2) Normalnormal (placing the memory pages starting from local DRAM). In this larger-scale case with more severe memory thrashing, as that prioritized page allocation to fast memory and triggered page migration only under memory pressure. As shown in Figure 16, Nomad led to lower performance compared to Memtis due to massive page migration operations but still outperformed TPPoutperformed TPP due to its graceful performance degradation during thrashing but fell short of matching Memtis’s performance. The initial placement strategy did not substantially affect the results and performance under different placements eventually converged.
Figure 12: Performance comparisons with of PageRank between non-migration, TPP, Memtis, and Nomad. Performance is normalized to the approach with the lowest speed.
@@
Graph-based computation. We used PageRank [pagerankwiki], an application used to rank web pages. It involves iterative computations to determine the rank of each page, based on the link structures of the entire web. As the size of the dataset increases, the computational complexity also increases, making it both memory-intensive and compute-intensive. We used a benchmark suite [pagerank] to generate a synthetic uniform-random graph comprising $2^{26}$ vertices, each with an average of 20 edges. The RSS in this experiment was 22 GB, indicating that the memory pages were distributed at both the local DRAM and remote CXL memory or PM.
@@
Figure 12 illustrates that there was negligible variance in performance between scenarios with page migrations (using Nomad and TPP) and without page migrations (no migration). The results suggest that: 1) For non-latency-sensitive applications, such as PageRank, using CXL memory can significantly expand the local DRAM capacity without adversely impacting application-level performance. 2) In such scenarios, page migration appears to be unnecessary. These findings also reveal that the overhead associated with Nomad’s page migration minimally influences PageRank’s performance. Additionally, it was observed that among all evaluated scenarios, Memtis exhibited the least efficient performance.
Figure 16 shows the case when we scaled the RSS to a very large scale on Platform platforms C & D. When the PageRank program started, it first used up to 100GB memory, then its RSS size dropped to 45GB to 50GB. Nomad achieved 2x the performance of TPP (both platforms) and slightly better than Memtis (Plactform platform C), due to more frequent page migrations – the local DRAM (16 GB) was not large enough to accommodate the WSS in this case.
Figure 13: Performance comparisons with of Liblinear between non-migration, TPP, Memtis, and Nomad. Performance is normalized to the approach with the lowest speed.
Figure 14: Redis (large RSS).
Figure 15: Page ranking (large RSS).
Figure 16: Liblinear (large RSS).
@@
Machine learning. Our final evaluation of Nomad involved using the machine learning library Liblinear [liblinear], known for its large-scale linear classification capabilities. We executed Liblinear with an L1 regularized logistic regression workload , using 16 threads with an RSS of 10 GB. Prior to each execution, we used our tool to demote all memory pages associated with the Liblinear workload to the slower memory tier.
Figure 13 demonstrate demonstrates that both Nomad and TPP significantly outperformed “no migration” and Memtis across all platforms, with performance improvement ranging from 20% to 150%. This result further illustrates that when the WSS is smaller than the local DRAM, Nomad and TPP can substantially enhance application performance by timely migrating hot pages to the faster memory tier. Figure 16 shows that with a much larger model and RSS when running Liblinear, NOMAD consistently achieved high performance across all cases. In contrast, TPP’s performance significantly declined, likely due to inefficiency issues, as frequent, high bursts in kernel CPU time were observed during TPP execution.
@@
Our observations of application evaluation results across four different platforms reveal that application-level performance is influenced by the distinct characteristics of each platform . For example, Platform A represents a standard off-the-shelf machine, while Platform B is an engineering sample with undisclosed features that may affect various applications differently. This suggests that other significant performance-related factors could potentially overshadow the effects of page migrations.
Unlike micro-benchmarks, the success rate in
Workload type Success : Aborted
Liblinear (large RSS) on platform C 1:1.9
Liblinear (large RSS) on platform D 2.6:1
Redis (large RSS) on platform C 153:1
Redis (large RSS) on platform D 278.2:1
Table 4: The Success rate of transactional migration.
Migration success rate. As stated in Section 3.1, Nomad’s page migrations varied across different applications, as shown in transactional page migration may be aborted due to updates to the migrating page, resulting in wasted memory bandwidth and CPU cycles. Subsequent retries could also fail. A low success rate could negatively affect application performance. Table 4 . shows Nomad’s migration success rate for Liblinear and Redis on platforms C and D. We chose a large RSS for both applications and ensured there were sufficient cross-tier migrations. We observed a low success rate for Liblinear , while Redis had a very high success rate. Interestingly, this contrasted with Nomad’s performance : it was excellent with Liblinear but poor with Redis both under the large RSSconditionwith large RSS. This suggests that a high success rate in page migrations does not necessarily equate lead to high performance. A low success rate can indicate indicates that the pages being migrated by Nomad are also being modified by other processes, implying their “hotness”. Timely migration of such pages can benefit ongoing and future accessaccesses.
Summary. The results from micro-benchmarks and applications indicate that when the WSS was smaller than the performance tier, Nomad enabled workloads to maintain higher performance than Memtis through asynchronous, transactional page migrations. However, when the WSS was comparable to or exceeded the performance tier capacity, leading to memory thrashing, the page-fault-based migration in Nomad became detrimental to workload performance, underperforming Memtis in write operations. Notably, Nomad’s page shadowing feature preserved the efficiency of read operations even under severe memory thrashing, often maintaining comparable or superior performance to Memtis. In all test scenarios, Nomad significantly outperformed the state-of-the-art page-fault-based migration approach, TPP.
The evaluation results across four different platforms reveal the following observations: First, Nomad generally performed better on platform D, which was equipped with faster and larger CXL memory, compared to other platforms. Additionally, the reduced performance gap between fast and slow memory on platform D allowed Nomad to achieve greater performance gains than TPP, as the performance overhead from TPP’s synchronous page migration was more pronounced. Second, while platforms A and B showed similar behavior in micro-benchmarks, their application-level performance varied (slightly) across different applications, suggesting that specific CPU features (differing between the off-the-shelf Intel Sapphire Rapids CPU for platform A and the engineering sample for platform B) may affect the performance of page migration under more realistic workloads.
@@ 5 Discussions and Future Work
The key insight from Nomad’s evaluation is that page migration, especially under memory pressure, has a detrimental impact on overall application performance. While Nomad achieved graceful performance degradation and much higher performance than TPP, an approach based on synchronous page migration, its performance is sub-optimal compared to that without page migration. When the program’s working set exceeds the capacity of the fast tier, the most effective strategy is to access pages directly from their initial placement, completely disabling page migration. It is straightforward to detect memory thrashing, e.g., frequent and equal number of page demotions and promotions, and disable page migrations. However, estimating the working set size to resume page migration becomes challenging, as the working set now spans multiple tiers. It requires global memory tracking, which could be prohibitively expensive, to identify the hot data set that can potentially be migrated to the fast tier. We plan to extend Nomad to unilaterally throttle page promotions and monitor page demotions to effectively manage memory pressure on the fast tier. Note that this would require the development of a new page migration policy, which is orthogonal to the Nomad page migration mechanisms proposed in this work.
Impact of Platform Characteristics: There exist difficult tradeoffs between page fault-based access tracking, such as TPP and Nomad, and hardware performance counter-based memory access sampling like Memtis. While page fault-based tracking effectively captures access recency, it can be potentially expensive and on the critical path of program execution. In comparison, hardware-based access sampling is off the critical path and captures access frequency. However, it is not responsive to workload changes and its accuracy relies on the sampling rate. One advantage of Nomad is that it is a page fault-based migration approach that is asynchronous and off the critical path. A potential future work is integrating Nomad with hardware-based, access frequency tracking, such as Memtis, to enhance the current migration policy.
@@ 6 Conclusion
This paper introduces non-exclusive memory tiering as an alternative to the common exclusive memory tiering strategy, where each page is confined to either fast or slow memory. The proposed approach, implemented in Nomad, leverages transactional page migration and page shadowing to enhance page management in Linux. Unlike traditional page migration, Nomad ensures asynchronous migration and retains shadow copies of recently promoted pages. Through comprehensive evaluations, Nomad demonstrates up to 6x performance improvement over existing methods, addressing critical performance degradation issues in exclusive memory tiering, especially under memory pressure. The paper calls for further research in tiered memory-aware memory allocation.
7 Acknowledgments
We thank our shepherd, Sudarsun Kannan, and the anonymous reviewers for their constructive feedback. This work was supported in part by NSF grants CCF-1845706, CNS-2415774, CCF-2415473, and the gift and equipment from Intel Labs and Micron Technology.
Appendix A Artifact Appendix
Abstract
The artifact contains the source code of Nomad, TPP, and Memtis for reproducing the results and graphs presented in the paper. The code works on platforms with Persistent Memory, Intel Agilex CXL memory, and/or Micron CXL memory. To facilitate the reproduction, we have provided a collection of scripts for compiling and installing these approaches, executing the experiments, collecting logs, and creating graphs. More details are available in the "README.md" file.
Scope
This artifact demonstrates Nomad’s strengths and weaknesses over TPP and Memtis across various scenarios and platforms, as elaborated in the Evaluation section.
It is open-source and can be used for further research, development, or other purposes by the community.
Contents
Nomad, TPP and Memtis implementation. We provide two separate patches to enable Nomad and TPP to work with the upstream kernel version v5.13-rc6. In particular, the TPP patch comes from the Linux community email discussions. Memtis, on the other hand, is directly incorporated from its original artifact, with a few minor bugs fixed.
Documentation The "Reproducing Paper Results" section of "README.md" provides a step-by-step guide for reproducing the results in the paper. This guide includes instructions for compiling the three implementations (i.e., Nomad, TPP, and Memtis), running the experiments, and generating the graphs as presented in the paper.
Hosting
Artifact link: https://github.com/lingfenghsiang/Nomad
Artifact license: GNU GPL V3.0
Artifact version tag: v0.0
Requirements
To reproduce the results in the paper, the system under test requires one NUMA node with a CPU and another CPU-less NUMA node. If the system has more NUMA nodes, the operating system might encounter unexpected errors. Additionally, Memtis is only fully functional on platforms with Optane Persistent Memory. More details are included in the "Prerequisites" section of "README.md".